OpenStudy (anonymous):

Find the exact value of the remaining trigonometric functions of theta: sec(theta)=4 , cot>0.

OpenStudy (anonymous):

ok, first of all \[\sec \theta = 1/\cos \theta \]

OpenStudy (anonymous):

Since sec (theta) = 4 we can say 1/cos (theta) = 4

OpenStudy (anonymous):

true

OpenStudy (anonymous):

\[1/\cos \theta = 4\] multiply both sides by cos (theta) \[4 \cos \theta = 1 \]

OpenStudy (anonymous):

cos = 1/4 then? which makes x=1 and radius of the unit circle 4?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok how do I find sin(theta) from there?

OpenStudy (anonymous):

To get sin (theta) you need to find y. Use the pythagorean theorem a^2 + b^2 = c^2, (a=x, b=y, and c=r) and since both the sec and cot are positive, the angle is in the first quadrant, so y is also going to be positive.

OpenStudy (anonymous):

1+ y^2 = 16 y^2 = 15 \[y = \sqrt{15}\]

OpenStudy (anonymous):

i got root(15) from that for y. so sin would be root(15) or since sine is y over radius would it be root(15)/4?

OpenStudy (anonymous):

since sin is y/r you need \[\sqrt{15}/4\]

OpenStudy (anonymous):

Alright I think I have it from here. thanks so much for your help.

OpenStudy (anonymous):

yup.